# Data Taggants: Dataset Ownership Verification via Harmless Targeted Data Poisoning

### v1

@@ 4.3 Robustness
For a fixed model and training algorithm, Table 1 and 2 already show the robustness of our method to a complete model retraining. But for our method to be practical, we need to ensure it will be robust when Bob’s model or training algorithm is different from Alice’s.
@@
Table 4 (and Table 8 in Appendix A.1) shows how even when Alice does not know Bob’s data augmentations, she can still successfully trace her dataset usage through Bob’s training, even though with much higher confidence when the same data augmentation is used by both Alice and Bob.
@@
Bob’s data aug. Alice’s data aug. Validation accuracy Top-$k$ keys accuracy
@@
Robustness “stress test”: different model architectures and data augmentation.
@@
We finally explore the robustness of data taggants in the most difficult setting, where Alice and Bob use different model architecture or sizes (see Table 11 in Appendix), as well as different data augmentations. We consider two families of models (DeIT and ResNet) of various sizes. We measure intra-family transferability by validating the protected dataset generated from each model onto each other model of the same family (DeIT and ResNet). We also measure inter-family transferability by crafting a protected dataset with ResNet-18 and DeIT-tiny models and validating it on all the models of the other family (resp. DeIT and ResNet). The results are shown in Table 5. Overall, we see good intra-family transferability, with larger DeIT models being more sensitive to taggants than smaller DeIT models. Interestingly, the trend is reversed for ResNets, with smaller ResNets being more sensitive. Across architectures, DeIT-tiny to ResNets or ResNets to DeIT-tiny, the results are less conclusive even though the top-10 accuracy is still $>0$. All in all, these results suggest that the taggants are robust even in this worst-case scenario, but getting high-confidence when Alice and Bob use different architectures and data augmentations is an interesting avenue for future work.
@@
Bob’s model Alice’s model Validation accuracy Top-$k$ keys accuracy
@@
ResNet-18 $67.0\pm 0.8$ $0.0\pm 0$ 0.0 ( 0.0) $10.0\pm 0.0$ ( -1.8)
DeIT-small DeIT-medium $64.5\pm 0.6$ $0.0\pm 0$ 0.0 ( 0.0) $22.5\pm 9.6$ ( -7.8)
@@
Limitations and future work.
This article presents a dataset ownership verification approach that marks a trained model. Current results show a slight degradation in validation accuracy compared to training on the clean dataset, which future works should try to reduce. While we show the robustness of our method through different transferability results against different model architectures and training recipe, further experiments would be required to evaluate the effectiveness of our approach when Bob trains his model on a subset or a superset of our protected dataset. Particularly, future work should study the efficiency of our method when Bob trains his model on an aggregate of different datasets containing data taggants.
5 Conclusion
We proposed data taggants, a new approach to dataset ownership verification and designed mostly for image classification datasets. Data taggants are hidden in a dataset through gradient matching, in order to mark models trained on them. Our approach shows promising results, with very high detection confidence of dishonest models in favorable settings, without affecting classification performance.
We highlight the difficulty of assessing stealthiness, a challenge that is likely faced by other related domains such as data poisoning and adversarial examples. Through extensive robustness tests, we also show the relevance of data taggants in difficult settings, and identify promising avenues for future work such as designing taggants that robustly mark models across large variations in training settings.
Appendix A Appendix
A.1 Ablations
Visual inspections.
We show, in Figure 5, randomly chosen samples of data taggants generated with and without perceptual loss.
Comparison of the perceptual loss with weight decay.
We ensure that the gain in PSNR offered by the perceptual loss is not simply due to it reducing the perturbation amplitudes. Figure 6 compare data taggants generated with perceptual loss, with weight decay on the perturbation (replacing the perceptual loss term by the norm 2 of the perturbation $\|\delta_{i}\|^{2}_{2}$), and with their vanilla counterpart. It shows that for a similar PSNR, the perceptual loss plays an important role in hiding the signature.
(a) vanilla
$\text{PSNR}=27.6\text{dB}$.
(b) with perceptual loss
$\text{PSNR}=30.6\text{dB}$.
(c) with weight decay
$\text{PSNR}=30.7\text{dB}$.
Figure 6: Comparison of signatures (top – amplified ×$\mathbf{\times 10}$) and data taggants (bottom) generated without discretion mechanism (left), with perceptual loss (center) and with weight decay (right).
Clean performances.
Data aug. Model Validation accuracy
SA DeIT-small $56.1\pm 0.5$
3A DeIT-medium $67.3\pm 1.1$
DeIT-small $64.2\pm 0.4$
DeIT-tiny $53.6\pm 0.4$
ResNet-50 $77.9\pm 0.0$
ResNet-34 $74.1\pm 0.0$
ResNet-18 $69.6\pm 0.0$
Baselines.
Similarly to our experiments, we report in Table 7, the validation accuracies and keys accuracies for our method and different baselines for keys sampled randomly or among the test set.
Key source Method Budget $B$ Validation accuracy Top-$k$ keys accuracy PSNR
$k=1$ $\log_{10}p$ $k=10$ $\log_{10}p$
None Clean training $0.0\%$ $64.2\pm 0.4$
Test data Naive Canary (Label Flipping) $0.001\%$ $63.8\pm 0.5$ $0.0\pm 0$ 0.0 (0.0) $0.0\pm 0$ 0.0 (0.0) -
$0.01\%$ $64.2\pm 0.7$ $57.5\pm 5.0$ (-54.0) $97.5\pm 5.0$ (-71.0) -
$0.1\%$ $63.6\pm 1.0$ $100.0\pm 0.0$ (-113.4) $100.0\pm 0.0$ (-74.0) -
Transparency $0.001\%$ $64.2\pm 1.1$ $0.0\pm 0$ 0.0 (0.0) $0.0\pm 0$ 0.0 (0.0) $20.0$
$0.01\%$ $63.2\pm 0.5$ $0.0\pm 0$ 0.0 (0.0) $0.0\pm 0$ 0.0 (0.0) $20.0$
$0.1\%$ $63.7\pm 0.6$ $0.0\pm 0$ 0.0 (0.0) $0.0\pm 0$ 0.0 (0.0) $20.0$
Data taggants $0.001\%$ $63.7\pm 0.4$ $0.0\pm 0$ 0.0 (0.0) $7.5\pm 5.0$ (-1.1) $27.5$
$0.01\%$ $63.4\pm 1.0$ $0.0\pm 0$ 0.0 (0.0) $0.0\pm 0$ 0.0 (0.0) $27.9$
$0.1\%$ $63.9\pm 0.6$ $0.0\pm 0$ 0.0 (0.0) $27.5\pm 5.0$ (-10.4) $28.4$
Random data Naive Canary $0.001\%$ $63.4\pm 0.7$ $7.5\pm 5.0$ (-3.3) $10.0\pm 0.0$ ( -1.8) -
$0.01\%$ $64.2\pm 0.3$ $15.0\pm 10.0$ (-9.2) $100.0\pm 0.0$ (-74.0) -
$0.1\%$ $63.8\pm 1.1$ $85.0\pm 19.1$ (-91.7) $100.0\pm 0.0$ (-74.0) -
$0.01\%$ $63.4\pm 0.7$ $7.5\pm 5.0$ (-3.3) $10.0\pm 0.0$ ( -1.8) $20.0$
$0.1\%$ $63.6\pm 0.6$ $10.0\pm 0.0$ (-4.9) $55.0\pm 5.8$ (-29.7) $20.0$
Data taggants (Our method) $0.001\%$ $63.7\pm 0.9$ $0.0\pm 0$ 0.0 (0.0) $0.0\pm 0$ 0.0 (0.0) $26.6$
$0.01\%$ $63.6\pm 0.9$ $0.0\pm 0$ 0.0 (0.0) $5.0\pm 5.8$ (-0.5) $27.3$
$0.1\%$ $64.2\pm 0.6$ $10.0\pm 0.0$ (-4.9) $87.5\pm 5.0$ (-59.6) $27.9$
(a) Key image $x^{(key)}$
(b) Transparency
(c) Data taggant
(d) Naive canary
Robustness to data augmentation changes.
Alice’s data aug. Bob’s data aug. Key source Validation accuracy Top-$k$ keys accuracy
$k=1$ $\log_{10}p$ $k=10$ $\log_{10}p$
Test data $56.2\pm 0.6$ $15.0\pm 17.3$ (-9.9) $60.0\pm 14.1$ (-34.2)
Test data $63.9\pm 1.1$ $0.0\pm 0$ 0.0 ( 0.0) $2.5\pm 5.0$ ( -0.1)
3A SA Random $56.1\pm 0.3$ $1.7\pm 4.1$ ( 0.0) $15.0\pm 8.4$ ( -1.8)
Test data $56.1\pm 0.4$ $0.0\pm 0$ 0.0 ( 0.0) $\mathbf{16.0\pm 5.5}$ ( -3.9)
Test data $63.9\pm 0.6$ $0.0\pm 0$ 0.0 ( 0.0) $27.5\pm 5.0$ (-10.4)
@@ A.1 Ablations
Key source Validation accuracy Top-$k$ keys accuracy
@@
Test data $79.4\pm 0.3$ $0.0\pm 0$ 0.0 ( 0.0) $0.0\pm 0$ 0.0 ( 0.0)
A.2 Experimental details
@@
We estimate the total computation time for this project (including preliminary runs) at around 150k GPU-hours.
@@
We present here the list of models used and their size and number of parameters

### v2

@@ 4.3 Robustness
For a fixed model and training algorithm, Table 2 and 3 already show the robustness of our method to a complete model retraining. But for our method to be practical, we need to ensure it will be robust when Bob’s model or training algorithm is different from Alice’s.
@@
Table 4 (and Table 8 in Appendix A.2) shows how even when Alice does not know Bob’s data augmentations, she can still successfully detect Bob’s model, even though we gain much higher confidence when the same data augmentation is used by both Alice and Bob. Also, since the 3A data augmentation uses mixup (Zhang, 2017) and cutmix (Yun et al., 2019), these results also demonstrate the robustness of our approach to Borgnia et al. (2021) which showed that mixup and cutmix can be used to defend against data poisoning attacks.
@@
Bob’s
data aug.
Alice’s
data aug.
Validation
accuracy
Top-$k$ keys accuracy
@@
“Stress test”: different model architectures and data augmentation.
@@
We finally explore the robustness of data taggants in the most difficult setting, where Alice and Bob use different data augmentations as well as model architecture or sizes (see Table 13 in Appendix). We consider two families of models (DeIT and ResNet) of various sizes. We measure intra-family transferability by validating the protected dataset generated from each model onto each other model of the same family (DeIT and ResNet). We also measure inter-family transferability by crafting a protected dataset with ResNet-18 and DeIT-tiny models and validating it on all the models of the other family (resp. DeIT and ResNet). The results are shown in Table 9 in Appendix. Overall, we see good intra-family transferability, with larger DeIT models being more sensitive to taggants than smaller DeIT models. Interestingly, the trend is reversed for ResNets, with smaller ResNets being more sensitive. Across architectures, DeIT-tiny to ResNets or ResNets to DeIT-tiny, the results are less conclusive even though the top-10 accuracy is still >0>0. All in all, these results suggest that the taggants are robust even in this worst-case scenario.
Dataset change.
We explore the case where Bob trains his model on a modified version of Alice’s dataset. We consider two cases: (1) Bob trains on multiple datasets (a superset of Alice’s dataset), and (2) Bob trains on a subset of Alice’s dataset. We present the results in Table 10 in Appendix and show data taggants are still effective in these cases.
5 Limitations and future work
While our non-backdoor dataset ownership verification approach shows good properties in terms of effectiveness, harmlessness, stealthiness, and robustness, limitations are to be observed. Current results show a negligible degradation in validation accuracy compared to training on the clean dataset, which future works should try to further reduce. While we show the robustness of our method through different transfer experiments against different model architectures and training recipe, by modifying by modifying them, Bob can still hurt the detection performances. Obtaining higher confidence when Alice and Bob use different architectures and data augmentations is an interesting avenue for future work.
6 Conclusion
We introduced data taggants, a new approach for dataset ownership verification and designed mostly for image classification datasets. Data taggants are hidden in a dataset through gradient matching, in order to mark models trained on them. Our approach shows promising results, with very high detection rate and confidence, and low false positive rate, without affecting models’ performance.
7 Ethics Statement
Our work is motivated by the need to ensure the integrity of machine learning models and the datasets they are trained on. While our approach to Dataset Ownership Verification (DOV) displays strong results, it is important to consider that such method can also fail. False positives can lead to misconceptions, particularly in such contexts.
8 Reproducibility Statement
The code implementing our method should be released upon publication. We provide all the necessary details to reproduce our experiments in the Section 4 and in the Appendix A.3.
Appendix A Appendix
A.1 Comparison with prior work
To visually clarify the position of our work w.r.t. prior work, we represent them in Table 5 to highlight the main dimensions of comparison: (i) what are the means of detection, (ii) whether the method relies on black-box access to the model, (iii) whether it is clean-label, (iv) imperceptibility, and (v) what are the theoretical guarantees. A yellow cross ✗ indicates slight disagreement (e.g. for the ‘black-box’ axis of analysis, for method that use logits as mean of detection, since it represents significatively more information than just sharing the predictions). Please note that this table is coarse and represent the best case scenario for cases where a variety of methods can be used (e.g. Backdoor watermarking).
mean of
detection
black-box clean-label imperceptible
theoretical
guarantees
Radioactive data
(Sablayrolles et al., 2020)
weights/
logits
on FPR
Dataset inference
(Maini, 2021)
logits/
pred.
N/A
on linear
models
Backdoor watermarking
(Li et al., 2022; 2023)
(Tang et al., 2023)
logits/
pred.
not
always
Data Isotopes
(Wenger et al., 2022)
logits/
top-$k$ pred.
Domain Watermark
Guo et al. (2023)
logits on risk
Data taggants (Ours)
top-$k$ pred. on FPR
Table 5: Comparison of DOV methods along dimensions representing different desirable properties.
Among the prior work, we chose to discard several of them from our pool of baselines. Domain watermark (Guo et al., 2023) has yet to share the implementation and complete details for generating their “hardly-generalized domains” which limits us in reproducing their method. It is comparable to Data taggants but targets “hard” samples from their hardly-generalized domains that are deemed unlikely to be well classified if not trained on the protected data. On the other hand, we provide theoretical guarantees that the keys are indeed unlikely to be well classified by a benign model. Dataset inference (Maini, 2021) in black-box requires a very large amount of queries (a few hundred per data point to verify), which makes it impractical. Radioactive data (Sablayrolles et al., 2020) in black-box can only be done via distillation of the model to evaluate, which also require a large number of queries to the model. Data taggants, on the other hand, only requires $K$ queries ($K=10$ in our experiments). We thus kept as baselines and compared against Backdoor watermarking and Data isotopes (Wenger et al., 2022). For Backdoor watermarking, we decided to leverage two approaches: a poisoned-label approach, BadNets (Gu et al., 2019), for its simplicity, and a clean-label approach, Sleeper Agent (Souri et al., 2021). The choice for this methods comes from the Domain Watermark (Guo et al., 2023) paper showing in Table 1 that Sleeper Agent performs better than the other tested clean-label backdoor watermarking approaches on Tiny-ImageNet.
A.2 Ablations
Visual inspections.
We show, in Figure 6, randomly chosen samples of data taggants generated with and without perceptual loss. The perceptual loss improves the stealthiness of the data taggants by making the perturbations less noticeable to the human eye. Some artifacts are still visible but could easily be overlooked by a human observer or misjudged as compression artifact or image grain.
Comparison of the perceptual loss with weight decay.
We ensure that the gain in PSNR offered by the perceptual loss is not simply due to it reducing the perturbation amplitudes. Figure 7 compare data taggants generated with perceptual loss, with weight decay on the perturbation (replacing the perceptual loss term by the norm 2 of the perturbation $\|\delta_{i}\|^{2}_{2}$), and with their vanilla counterpart. It shows that for a similar PSNR, the weight decay version is much more noticeable than its perceptual loss counterpart. The perceptual loss hence plays an important role in hiding the signature.
(a) vanilla
$\text{PSNR}=27.6\text{dB}$.
(b) with perceptual loss
$\text{PSNR}=30.6\text{dB}$.
(c) with weight decay
$\text{PSNR}=30.7\text{dB}$.
Figure 7: Comparison of signatures (top – amplified $\mathbf{\times 10}$) and data taggants (bottom) generated without discretion mechanism (left), with perceptual loss (center) and with weight decay (right).
Clean performances.
Data aug. Model Validation accuracy
SA DeIT-small $56.1\pm 0.5$
3A DeIT-medium $67.3\pm 1.1$
DeIT-small $64.2\pm 0.4$
DeIT-tiny $53.6\pm 0.4$
ResNet-50 $77.9\pm 0.0$
ResNet-34 $74.1\pm 0.0$
ResNet-18 $69.6\pm 0.0$
Baselines.
Similarly to our experiments, we report in Table 7, the validation accuracies and keys accuracies for our method and different baselines for keys uniformly sampled in the pixel space or from the test set. Figure 8 shows how Backdoor Watermarking-based detection (Li et al., 2023) performs with two backdoor attacks: Sleeper Agent (Souri et al., 2021) and BadNets (Gu et al., 2019). As Sleeper Agent displays a high $p$-value for both the watermarked model and the benign model, any relevant level of significance (below $0.05$) yield low detection rate (True Positive Rate) and low false detection rate (False Positive Rate). On the other hand, as we lower the hyperparameter of the tested hypothesis, the $p$-value for associated with the benign model is lower than that of the watermarked model, which should lead to false detection rate higher than the detection rate. For BadNets, the $p$-value obtained from the t-test is lower for benign model than watermarked models as the threshold decreases, leading to a potentially high false detection rate.
(a) Sleeper Agent
(b) BadNets
Figure 8: $p$-values computed for the Backdoor Watermarking-based detection (Li et al., 2023) as a function of the margin, the threshold of their null hypothesis.
Key
source
Method
Budget
$B$
Validation
accuracy
Top-$k$ keys accuracy PSNR
$k=1$ $\log_{10}p$ $k=10$ $\log_{10}p$
None Clean training $0.0\%$ $64.2\pm 0.4$ - - - - -
Test
data
Naive Canary
(Label Flipping)
$0.001\%$ $63.8\pm 0.5$ $0.0\pm 0.0$ (0.0) $0.0\pm 0.0$ (0.0) -
$0.01\%$ $64.2\pm 0.7$ $57.5\pm 5.0$ (-54.0) $97.5\pm 5.0$ (-71.0) -
$0.1\%$ $63.6\pm 1.0$ $100.0\pm 0.0$ (-113.4) $100.0\pm 0.0$ (-74.0) -
Transparency $0.001\%$ $64.2\pm 1.1$ $0.0\pm 0.0$ (0.0) $0.0\pm 0.0$ (0.0) $20.0$
$0.01\%$ $63.2\pm 0.5$ $0.0\pm 0.0$ (0.0) $0.0\pm 0.0$ (0.0) $20.0$
$0.1\%$ $63.7\pm 0.6$ $0.0\pm 0.0$ (0.0) $0.0\pm 0.0$ (0.0) $20.0$
Data taggants
$0.001\%$ $63.7\pm 0.4$ $0.0\pm 0.0$ (0.0) $7.5\pm 5.0$ (-1.1) $27.5$
$0.01\%$ $63.4\pm 1.0$ $0.0\pm 0.0$ (0.0) $0.0\pm 0.0$ (0.0) $27.9$
$0.1\%$ $63.9\pm 0.6$ $0.0\pm 0.0$ (0.0) $27.5\pm 5.0$ (-10.4) $28.4$
Random
data
$0.01\%$ $64.2\pm 0.3$ $15.0\pm 10.0$ (-9.2) $100.0\pm 0.0$ (-74.0) -
$0.1\%$ $63.8\pm 1.1$ $85.0\pm 19.1$ (-91.7) $100.0\pm 0.0$ (-74.0) -
$0.01\%$ $63.4\pm 0.7$ $7.5\pm 5.0$ (-3.3) $10.0\pm 0.0$ ( -1.8) $20.0$
$0.1\%$ $63.6\pm 0.6$ $10.0\pm 0.0$ (-4.9) $55.0\pm 5.8$ (-29.7) $20.0$
Data taggants
(Our method)
$0.001\%$ $63.7\pm 0.9$ $0.0\pm 0.0$ (0.0) $0.0\pm 0.0$ (0.0) $26.6$
$0.01\%$ $63.6\pm 0.9$ $0.0\pm 0.0$ (0.0) $5.0\pm 5.8$ (-0.5) $27.3$
$0.1\%$ $64.2\pm 0.6$ $10.0\pm 0.0$ (-4.9) $87.5\pm 5.0$ (-59.6) $27.9$
(a) Key image $x^{(key)}$
(b) Transparency
(c) Data taggant
(d) Naive canary
Robustness to data augmentation changes.
We show in Table 8 that our method is robust to data augmentation changes. We also confirm that uniformly sampling keys in the pixel space to be out-of-distribution is a better strategy than using test data as keys.
Alice’s
data aug.
Bob’s
data aug.
Key
source
Validation
accuracy
Top-$k$ keys accuracy
$k=1$ $\log_{10}p$ $k=10$ $\log_{10}p$
Test data $56.2\pm 0.6$ $15.0\pm 17.3$ (-9.9) $60.0\pm 14.1$ (-34.2)
Test data $63.9\pm 1.1$ $0.0\pm 0.0$ ( 0.0) $2.5\pm 5.0$ ( -0.1)
3A SA Random $56.1\pm 0.3$ $1.7\pm 4.1$ ( 0.0) $15.0\pm 8.4$ ( -1.8)
Test data $56.1\pm 0.4$ $0.0\pm 0.0$ ( 0.0) $\mathbf{16.0\pm 5.5}$ ( -3.9)
Test data $63.9\pm 0.6$ $0.0\pm 0.0$ ( 0.0) $27.5\pm 5.0$ (-10.4)
Robustness to stress test.
On top of the data augmentation changes, we show in Table 9 that our method displays some robustness to changes in the model architecture too.
@@ A.2 Ablations
Bob’s
model
Alice’s
model
Validation
accuracy
Top-$k$ keys accuracy
@@
ResNet-18 $67.0\pm 0.8$ $0.0\pm 0.0$ ( 0.0) $10.0\pm 0.0$ ( -1.8)
DeIT-small DeIT-medium $64.5\pm 0.6$ $0.0\pm 0.0$ ( 0.0) $22.5\pm 9.6$ ( -7.8)
@@
Dataset change.
We consider two cases:
(1) Bob trains on a superset of Alice’s dataset. In our experiments, we make Alice use only half the classes of ImageNet1k and Bob trains on the whole dataset.
(2) Bob subsamples Alice’s dataset. We make Alice use only half the classes of ImageNet1k. Bob will then remove 20% of Alice’s samples from each class and add in the classes that do not belong to Alice.
Results are showed in Table 10 and show that in both cases, our method is still able to detect Bob’s model strong reaction to the keys.
Table 10: Results of our method when Bob trains his model on a modified version of Alice’s dataset.
Case Validation accuracy Top-10 key accuracy $\log_{10}p$
(1) $64.5\pm 0.6$ $62.5\pm 15.0$ (36.3)(-36.3)
(2) $58.6\pm 0.7$ $72.5\pm 9.6$ (44.9)(-44.9)
@@
Key
source
Validation
accuracy
Top-$k$ keys accuracy
@@
Test data $79.4\pm 0.3$ $0.0\pm 0.0$ ( 0.0) $0.0\pm 0.0$ ( 0.0)
A.3 Experimental details
@@
We present in Table 13 the list of models used and their size and number of parameters. Table 13 details the data augmentations used in our experiments. Finally, Table 14 presents the training recipe used for our experiments.